Molybdenum forms body-centered cubic crystals and at 20 degree celsius the density is 10.3g/ml. Calculate the distance between the centers of the nearest molybdenum atoms.


1
Expert's answer
2020-08-27T05:05:02-0400

Density = Z × M / N × a3.


Z is effective number of atoms for Molybdenum (Mo). = 2 .


M is Molar mass of Mo = 95.95 g/ mol.


N = 6.02 × 1023.


a is edgelength in Cm .


Putting all Value


d = 10.3 g / cm^3.


M = 95.95 g/ mol.


N = 6.02 × 1023.


Then find "a" from above formula .


a = 3.16 × 10-8 Cm.


This is edge length and minimum dustance betweeb center of Mo .